# In-Context Editing: Learning Knowledge from Self-Induced Distributions

### v3

@@ 1 Introduction
Knowledge editing [55] has emerged as a research area that addresses the challenge of efficiently updating LLM outputs in response to specific queries. It focuses on modifying particular pieces of knowledge in a language model $\mathcal{M}_{\theta}$ using query-response pairs $\{(\mathbf{q}_{i},\mathbf{x}_{i}^{*})\}_{i=1}^{N}$. For instance, given the query “The president of the US is”, a model trained on outdated data might respond “Donald Trump”, while the desired up-to-date response would be “Joe Biden”.
@@ 2.1 Knowledge Editing: Problem Setup
The objective of knowledge editing is to incorporate new facts into a language model $\mathcal{M}_{\theta}$ through query-response pairs $\{(\mathbf{q}_{i},\mathbf{x}_{i}^{*})\}_{i\in[1,N]}$. Here $\mathbf{q}$ is the query that triggers the retrieval of factual knowledge from $\mathcal{M}_{\theta}$, such as ”The president of the US is”, with $\mathbf{x}^{*}$ being the intended response after editing, e.g., ”Joe Biden”. This integration is typically done by maximizing the probability $p_{\theta}(\mathbf{x}^{*}|\mathbf{q})$. Conventionally, a knowledge editing algorithm is assessed across four key dimensions.
@@
Portability assesses how well the model generalizes the knowledge for rephrased or logically related queries within the edit scope $\mathcal{D}_{q}$. For the aforementioned example, a related query could be ”The first lady of the US is”, with the target being ”Jill Biden” instead of ”Melania Trump”.
@@ 2.2 Knowledge-Editing Approaches
Another line of work, which our method focuses on, involves editing the model’s weights to integrate new knowledge. Approaches include direct direct fine-tuning [59; 23; 34], meta-learning-driven approaches [43; 8; 30; 41], and targeted network editing [10; 7; 25; 5; 38; 52; 44; 20; 13; 15; 16; 1; 11; 53; 51; 26; 12; 46; 39; 37]. These methods aim to incorporate target knowledge and employ various techniques to ensure the locality of the edits. Techniques include constraining the gradient of parameters [59], adopting parameter-efficient approaches [54], and applying statistical constraints on the weights, with notable examples such as ROME [28] and MEMIT [29]. In this paper, we propose tuning the model towards a self-generated distribution instead of a one-hot target.
@@ 3 Methodology: Learning Knowledge From Contexts
Running Example: To illustrate our approach, suppose we want to update a language model to reflect that the current president of the United States is Joe Biden, whereas it currently outputs Donald Trump when queried with “The president of the US is”. Our goal is to update the model so that it outputs Joe Biden for this query, without affecting its performance on unrelated queries.
@@ 3.1 Vanilla Fine-Tuning
In our running example, this means we fine-tune the model to assign maximum probability to the sequence “Joe Biden” given the query “The president of the US is”. While this approach can effectively update the model’s response for a specific query, it has significant drawbacks. The use of a one-hot target distribution often leads to overfitting, causing the model to degrade, suffer from catastrophic forgetting, or even collapse, resulting in unnatural or repetitive outputs.
@@ 3.3 In-Context Tuning with Sampling
To address the ineffectiveness of the naive sampling approach in § 3.2, we introduce extra information that guides the model towards a new distribution that aligns with the target, while maintaining similarity to its original distribution. Specifically, we leverage the in-context learning capabilities of language models by prepending context prompts $\mathbf{c}$ to the queries $\mathbf{q}$, where $\mathbf{c}$ is the new knowledge to be learned. For our example, we can create a context such as “Joe Biden is the current president of the United States.”, and prepend this context to the query. This induces a new contextual distribution $p_{\theta_{0}}(\mathbf{x}|[\mathbf{c},\mathbf{q}])$ that incorporates the desired knowledge through the context, while keeping minimal changes to the model. We can define the loss function as:
@@ 3.4 Consistent In-Context Editing (ICE)
Optimizing the Combined Loss $\mathcal{L}$: To optimize the total loss as defined in Equation 6, we sample sequences $\mathbf{x}_{\mathbf{c}}$ from the model conditioned on $[\mathbf{c},\mathbf{q},\mathbf{x}^{*}]$ and maximize the likelihood of the combined sequence $[\mathbf{x}^{*},\mathbf{x}_{\mathbf{c}}]$. This process is equivalent to optimizing the combined loss $\mathcal{L}$. If the sampling is not conditioned on the target, we would be solely optimizing $L_{ICE}$. This approach is algorithmically convenient, and we provide a proof of this equivalence in § A.3. To prevent the model from drifting too far from the initial parameters (thus preserving unrelated knowledge), we employ gradient clipping techniques inspired by constrained fine-tuning methods [59]. The detailed algorithm is presented in Algorithm 1.
@@ A.3 Decomposing Consistent In-Context Editing
$=\mathcal{L}_{ft}+\mathcal{L}_{ice}\Bigl{(}[\mathbf{q},\mathbf{x}^{*}]\Bigr{)},$
@@ Appendix B Algorithm
Here we provide the algorithm of in-context editing (ICE) in Algorithm 1.

### v4

@@ 1 Introduction
Knowledge editing [55] has emerged as a research area that addresses the challenge of efficiently updating LLM outputs in response to specific queries. It focuses on modifying particular pieces of knowledge in a language model $\mathcal{M}_{\theta}$ using query-response pairs $\{(\mathbf{q}_{i},\mathbf{x}_{i}^{*})\}_{i=1}^{N}$. For instance, given the query “The president of the US is”, a model trained on outdated data might respond “Joe Biden”, while the up-to-date response would be “Donald Trump”.
@@ 2.1 Knowledge Editing: Problem Setup
The objective of knowledge editing is to incorporate new facts into a language model $\mathcal{M}_{\theta}$ through query-response pairs $\{(\mathbf{q}_{i},\mathbf{x}_{i}^{*})\}_{i\in[1,N]}$. Here $\mathbf{q}$ is the query that triggers the retrieval of factual knowledge from $\mathcal{M}_{\theta}$, such as “The president of the US is”, with $\mathbf{x}^{*}$ being the intended response after editing, e.g., “Donald Trump”. This integration is typically done by maximizing the probability $p_{\theta}(\mathbf{x}^{*}|\mathbf{q})$. Conventionally, a knowledge editing algorithm is assessed across four key dimensions.
@@
Portability assesses how well the model generalizes the knowledge for rephrased or logically related queries within the edit scope $\mathcal{D}_{q}$. For the aforementioned example, a related query could be ”The first lady of the US is”, with the target being ”Melania Trump” instead of ”Jill Biden”.
@@ 2.2 Knowledge-Editing Approaches
Another line of work, which our method focuses on, involves editing the model’s weights to integrate new knowledge. Approaches include direct fine-tuning [59; 23; 34], meta-learning-driven approaches [43; 8; 30; 41], and targeted network editing [10; 7; 25; 5; 38; 52; 44; 20; 13; 15; 16; 1; 11; 53; 51; 26; 12; 46; 39; 37]. These methods aim to incorporate target knowledge and employ various techniques to ensure the locality of the edits. Techniques include constraining the gradient of parameters [59], adopting parameter-efficient approaches [54], and applying statistical constraints on the weights, with notable examples such as ROME [28] and MEMIT [29]. In this paper, we propose tuning the model towards a self-generated distribution instead of a one-hot target.
@@ 3 Methodology: Learning Knowledge From Contexts
Running Example: To illustrate our approach, suppose we want to update a language model to reflect that the current president of the United States is Donald Trump, whereas it currently outputs Joe Biden when queried with “The president of the US is”. The model needs to be updated so that it outputs Donald Trump for this query, without affecting its performance on unrelated queries.
@@ 3.1 Vanilla Fine-Tuning
In our running example, this means we fine-tune the model to assign maximum probability to the sequence “Donald Trump” given the query “The president of the US is”. While this approach can effectively update the model’s response for a specific query, it has significant drawbacks. The use of a one-hot target distribution often leads to overfitting, causing the model to degrade, suffer from catastrophic forgetting, or even collapse, resulting in unnatural or repetitive outputs.
@@ 3.3 In-Context Tuning with Sampling
To address the ineffectiveness of the naive sampling approach in § 3.2, we introduce extra information that guides the model towards a new distribution that aligns with the target, while maintaining similarity to its original distribution. Specifically, we leverage the in-context learning capabilities of language models by prepending context prompts $\mathbf{c}$ to the queries $\mathbf{q}$, where $\mathbf{c}$ is the new knowledge to be learned. For our example, we can create a context such as “Donald Trump is the current president of the United States.”, and prepend this context to the query. This induces a new contextual distribution $p_{\theta_{0}}(\mathbf{x}|[\mathbf{c},\mathbf{q}])$ that incorporates the desired knowledge through the context, while keeping minimal changes to the model. We can define the loss function as:
@@ 3.4 Consistent In-Context Editing (ICE)
Optimizing the Combined Loss $\mathcal{L}$: To optimize the total loss as defined in Equation 6, we sample sequences $\mathbf{x}_{\mathbf{c}}$ from the model conditioned on $[\mathbf{c},\mathbf{q},\mathbf{x}^{*}]$ and maximize the likelihood of the combined sequence $[\mathbf{x}^{*},\mathbf{x}_{\mathbf{c}}]$. This process is equivalent to optimizing the combined loss $\mathcal{L}$. If the sampling is not conditioned on the target, we would be solely optimizing $L_{\text{ICE}}$. This approach is algorithmically convenient, and we provide a proof of this equivalence in § A.3. To prevent the model from drifting too far from the initial parameters (thus preserving unrelated knowledge), we employ gradient clipping techniques inspired by constrained fine-tuning methods [59]. The detailed algorithm is presented in Algorithm 1.
@@ A.3 Decomposing Consistent In-Context Editing
$=\mathcal{L}_{\text{FT}}+\mathcal{L}_{\text{ICE}}\Bigl{(}[\mathbf{q},\mathbf{x}^{*}]\Bigr{)},$
@@ Appendix B Algorithm
Here, we provide the algorithm of in-context editing (ICE) in Algorithm 1.